Circular No. 5936

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) IAUC 5906 discusses orbits that utilize the first postconjunction ob-

servations of nine of the nuclei. MPC 22931 contains orbits for two more nuclei, 16 and 19, for which preconjunction data could be linked with postconjunction data on two or more nights. MPC 23105–23107 include orbits for six more nuclei—2, 4, 8, 18, 20 and 21. However, the pre- to postconjunction linkage for nucleus 8 was incorrect. The correct linkage is shown on MPEC 1994-D02, which contains new orbits for nuclei 8 and 9. Much of the recent analysis depends on measurements by H. Weaver from images obtained with the Hubble Space Telescope (HST) during Jan. 24–27. These images, as well as the HST image of 1993 July 1, indicate that nuclei 7 and 8 are in fact double, the bright nucleus 7a having a faint companion 7b currently 1′′.2 to the north and the moderately bright nucleus 8b a somewhat fainter companion 8a currently 2′′.2 to the west. For orbit-determination purposes the original nuclei 7 and 8 are linked with the current 7a and 8b. The original nuclei 10 and 13 have disappeared. Nucleus 3 still appears to exist, but there are no other postconjunction measurements. The list of collision times with Jupiter given on IAUC 5906 can therefore be supplemented as follows: nucleus 21, July 16.8 UT; 20, 17.1; 19, 17.2; 18, 17.4; 16, 18.0; 9, 20.4; 8, 20.6; 4, 21.6; 2, 22.1.

NOVA CASSIOPEIAE 1993 M. Kidger, N. Devaney, K. Sahu and S. López, Instituto de Astrofísica

de Canarias, report that 1.4–2.4-μm CVF spectrophotometric and JHKL′ photometric monitoring with the 1.52-m Carlos Sánchez Telescope, Teide Observatory, shows major changes in the infrared flux distribution consistent with the formation of a dust shell, as predicted by Shore et al. (IAUC 5925) and coincident with the optical fading (IAUC 5934). The infrared spectrum has been regularly monitored since 1993 Dec. 16. The strong CO feature reported by Scott et al. (IAUC 5916) was still very prominent on Dec. 28, but it had faded considerably by 1994 Jan. 5, when the Brackett lines and Paα were extremely strong; these lines had weakened considerably by Feb. 5. On Feb. 15 the continuum was observed to be very red, consistent with blackbody emission from dust at ∼ 700 K. The spectral lines are now very weak or invisible. This spectral evolution is very similar to that of Nova Her 1991 = V838 Her. The infrared magnitudes H, K and L′ are seen to have brightened considerably since Feb. 2, viz.: Feb. 2.90 UT, J = 6.37, H = 6.04, K = 5.43, L′ = 3.77; 15.85, 6.64, 4.80, 3.13, 1.11; 16.84, 6.55, 4.68, 2.98, 0.96 (errors ±0.03).

1994 February 18 Brian G. Marsden